OpenStudy (anonymous):

A sample of sulfur hexafluoride gas occupies a volume of 9.10 L at 198°C. Assuming that the pressure remains constant, what temperature (in °C) is needed to reduce the volume to 2.50 L?

OpenStudy (zpupster):

the formula you need since pressure stays the same is: V1/T1=V2/T2 so V1=9.1L T1 = 198C V2=2.5L T2=??

OpenStudy (anonymous):

i posted this over estimating the problem, would you mind doing it out and seeing if you get 54.4 degrees C

OpenStudy (anonymous):

if my answers wrong then i may still need help

OpenStudy (anonymous):

otherwise thanks!

OpenStudy (zpupster):

good job!! thats what i got.